OpenStudy (anonymous):

Hannah can paint a room in 9 hours. Destiny can paint the same room in 15 hours. How long does it take for both Hannah and Destiny to paint the room it they are working together.

OpenStudy (anonymous):

Not 9+15 lol.

OpenStudy (anonymous):

no, it would be a reduced time

OpenStudy (anonymous):

I am thinking its 6, but thats wrong

OpenStudy (anonymous):

Here it is how much of the room can hannah paint in 1 hour? 1/20 how much of the room can destiny paint in 1 hour? 1/16 how much of the room can they paint working together in 1 hour? 1/20 + 1/16 = 1/20(4/4) + 1/16(5/5) = 4/80 + 5/80 = 9/80 all of the other work rates were the reciprocal of the hours needed for the full job. Same with this one. So hours for the full job would be 80/9, or 8.9 hours to complete the job together. 1/(hours for 1st person) + 1(hours for 2nd person) = 1/(hours together)

OpenStudy (anonymous):

I might've messed up with that, lemme seee

OpenStudy (anonymous):

you type really quick!!!!!!

OpenStudy (anonymous):

ok, but i got a decimal, which doesnt seem right

OpenStudy (anonymous):

Do I..?

OpenStudy (anonymous):

yes o.o

OpenStudy (anonymous):

Not too quick xD

OpenStudy (anonymous):

9/2= 4.5 15/2= 7.5 4.5+7.5=12 12 Hours

OpenStudy (anonymous):

So 12 is the answer

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

yes, you are probably right..but no, didnt take it

OpenStudy (anonymous):

?? Whatcha mean?

OpenStudy (anonymous):

Im sure the answer is 12 >.>

OpenStudy (anonymous):

its an online homework page, it didnt work :p

OpenStudy (anonymous):

ahgahga

OpenStudy (anonymous):

It's supposed to be reduced since it's 2 people which make it less time .

OpenStudy (anonymous):

Mhmm .-.